find the sum of the first 67 odd integers?
Answer by stanbon(75887)

find the sum of the first 67 odd integers?
--------
67th odd = 2n-1 = 2*67-1 = 133
The sequence: 1,3,5,7,.....,133
----
Sum = (67/2)(1+133) = 33.5*134 = 4489
---
Cheers,
Stan H.
================================